OpenStudy (anonymous):

For an angle Θ with the point (16, −30) on its terminating side, what is the value of cosine?

OpenStudy (anonymous):

Here are the ansers 16 −30 negative 15 over 17 8 over 17

OpenStudy (mathstudent55):



OpenStudy (anonymous):

how do i get cos

OpenStudy (mathstudent55):

What is the cosine defined as? Hint: do you know SOHCAHTOA?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what's the adjacent

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (mathstudent55):

\(\cos \theta = \dfrac{adj}{hyp} \) Remember this?

OpenStudy (mathstudent55):